OpenStudy (austinwst13):

What is the value of ? A. 4 B. 8 C. 22.63 D. 64 @rebeccaxhawaii

rebeccaxhawaii (rebeccaxhawaii):

of what?

OpenStudy (austinwst13):

3/512

rebeccaxhawaii (rebeccaxhawaii):

3/512=3/512

rebeccaxhawaii (rebeccaxhawaii):

please sent a screenshot.

rebeccaxhawaii (rebeccaxhawaii):

3/512 is different than \[\sqrt[3]{512}\]

rebeccaxhawaii (rebeccaxhawaii):

so _*_*_=512

OpenStudy (austinwst13):

ooohhhh

rebeccaxhawaii (rebeccaxhawaii):

remember how 3^3=3*3*3 well now we are working backwards we need to find what 3 numbers equal 512. Keep in mind it needs to be all the same number.

OpenStudy (austinwst13):

i know its not d

rebeccaxhawaii (rebeccaxhawaii):

keep going.

OpenStudy (austinwst13):

i cam up with 512.1169

rebeccaxhawaii (rebeccaxhawaii):

huge hint: http://openstudy.com/study#/updates/573029bce4b0c2a20a4c8b6b

OpenStudy (austinwst13):

is it b 8

rebeccaxhawaii (rebeccaxhawaii):

they have already given you the value which is 512. you cant just change the problem and come up with your own. also 22.63^3 would not be 512.1169 its 11589.205447 anyways yes b is correct bc 8*8*8=512

OpenStudy (austinwst13):

got it b

OpenStudy (austinwst13):

thx can i post my second question on here

rebeccaxhawaii (rebeccaxhawaii):

post a new thread please

OpenStudy (austinwst13):

k